Circular No. 8441

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOTICE TO SUBSCRIBERS Due to the wide public access to the Internet during the last decade,

the Central Bureau has experienced a significant, sustained loss in paying subscribers — at a gradual-but-steady rate that now has the total paying subscribers to these Circulars (printed plus electronic) down by about a quarter in the last five years. We have been reluctant to raise prices for fear of losing more paying subscribers, and indeed no such increase has occurred in the past decade. Readership of these Circulars is way up over the past decade, due both to the former free availability of them at the CBAT website and to over-liberal distribution within institutions (and groups) that have single (or no) paid subscriptions. However, subscription funds are still necessary for the work behind issuing these Circulars, so we have initiated two immediate actions: (1) to block (at least temporarily) free online access to these Circulars in favor of our paying subscribers and (2) to request that all groups and institutions with multiple readers take out ‘multiple’ subscriptions (a concept that is paralleled in subscriptions to other astronomical journals in the sense of ‘institutional rates’). We propose that groups and institutions pay their subscriptions as follows: 5–10 readers, twice the single rate; 11–20 readers, 3×; 21–30, 4×; 31–60, 5×; 61–90, 6×; 91–130, 7×; 131–180, 8×; 181–250, 9×; 251–320, 10×; 321–400, 11×; 401 or more, 12×. As has always been the case, astronomers (and groups) with limited funds can ask for aid in reducing the subscription costs.

VARIABLE STAR IN CEPHEUS R. Persson, Hindas, Sweden, reports that a USNO-B1.0 star located

at α = 22^{h}53^{m}33^{s}.28, δ = +62^{o}32′23′′.1 (equinox 2000.0; blue mag 20.2, red mag 15.7) was not visible on a red Palomar Sky Survey plate from 1953 October, but it is present on digitized Palomar Sky Survey plates taken on 1991 Sept. 2 (red, mag ∼ 15.7) and 1984 Aug. 28 (Quick-V). B. Reipurth, University of Hawaii (UH); and C. Aspin, Gemini Observatory, report that an R-band CCD image obtained at the UH 2.2-m telescope on 2004 Oct. 9 shows that this star (at R ∼ 17.3) is located at the apex of what appears to be an outflow cavity in a small dark cloud and is apparently identical with the bright infrared source 2MASS 22533325+6232235 (K = 8.20). A GMOS spectrum (resolution 6200), obtained on Nov. 6 at the Gemini-N telescope, shows deep and wide Hα and Na absorption lines on a red continuum. Thus, the variable appears to be a new highly reddened FU-Ori-type object that has erupted sometime during the past 50 years.

2004 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green